OpenStudy (sirm3d):

@lovekblue \[\lim_{x\rightarrow \infty} \frac{2^{x/2}}{3^{x/3}}\]

OpenStudy (sirm3d):

notice how l'hop will fail?

OpenStudy (anonymous):

yes, cuz i've tried it LOL

OpenStudy (anonymous):

since it is a 2^x , u will always get 2^x back

OpenStudy (anonymous):

2^(x/3)*

OpenStudy (sirm3d):

yep.

OpenStudy (sirm3d):

problems of this type, one-term-numerator and one-term denominator, combine them as a fraction raised to a common power.

OpenStudy (anonymous):

will remember that!

OpenStudy (anonymous):

can u show me some (inf)^ inf

OpenStudy (sirm3d):

inf^inf = inf

OpenStudy (anonymous):

I have one here, but it looks pretty mesy

OpenStudy (sirm3d):

#12?

OpenStudy (anonymous):

yes

OpenStudy (sirm3d):

that's 0^inf

OpenStudy (sirm3d):

after applying ln, ln y = lim x ln (3x/1+2x2])=(inf)(-inf)=-inf y = e^-inf = 1/e^inf = 1/inf = 0 no L'hop needed.

OpenStudy (sirm3d):

\[\Large r^\infty =\begin{cases}0, & 0<r<1\\ \infty, & r>1\end{cases}\]

OpenStudy (anonymous):

(inf)(-inf)=-inf < where did u get this from?

OpenStudy (sirm3d):

lim (3x) / (1+2x^2) = 0 lim ln [ (3x) / (1+2x^2) ] = -inf

OpenStudy (anonymous):

i still dont get how is 3x / 1 +2x^2 = 0 dont u sub inf in x?

OpenStudy (anonymous):

or is it because the degree of x is bigger in denominator?

OpenStudy (sirm3d):

ln (inf) = inf ln (0) = -inf e^inf = inf e^-inf = 0

OpenStudy (sirm3d):

BINGO!!! it's all about the degree.

OpenStudy (anonymous):

omg yes, i finally got something xD

OpenStudy (anonymous):

it's been a while since u last said bingo LOL

OpenStudy (sirm3d):

a lot of students go through L'Hops evaluating lim (x-->inf) f(x)/g(x) for algebraic functions f and g, without realizing that only the degree of f and g determine whether the answer is inf, some constant c, or 0.

OpenStudy (sirm3d):

consider lim (x->inf) [(7x^7 + 6x^6 - 1) / (9x^9-5)]

OpenStudy (sirm3d):

it would be a waste applying L'hop 7 times and conclude that lim = 0

OpenStudy (anonymous):

oh 0 !

OpenStudy (anonymous):

yea, my prof never taught us that either

OpenStudy (sirm3d):

just write it as (0x^9 + 7x^7 + 6x^6 - 1) / (9x^9 - 5) the answer is determined completely by the coefficients of the terms with the highest degree 0/9 = 0

OpenStudy (anonymous):

oh, so if it is 5 x^9 .... / 6 x^9 , then limit is 5/6

OpenStudy (sirm3d):

YES!

OpenStudy (anonymous):

do u think i'm ready for the exam >.> 26 more hrs ..

OpenStudy (sirm3d):

you need to double-check your solution, since you make mistakes plenty of times

OpenStudy (anonymous):

yea.. I should get at least 10% higher than what i have if i didnt make mistakes :/

OpenStudy (anonymous):

and for L'H, so do we first look for 0^ inf = 0 or inf ^ 0 = 1 first. If it is not there, then try to manipulate the function into inf/inf or 0/0?

OpenStudy (sirm3d):

right. L'Hop works only on inf/inf or 0/0

OpenStudy (anonymous):

and if the exponent is not 0 or inf, use ln to bring it down.

OpenStudy (sirm3d):

right. but make sure to write an equation before using ln.

OpenStudy (anonymous):

why..? i always ignore the y @@

OpenStudy (sirm3d):

yes.

OpenStudy (anonymous):

For lim x-> 0+ , e^ 1/x = inf ?

OpenStudy (anonymous):

isnt 1/0 u.d?

OpenStudy (sirm3d):

lim (3x/1+2x2)^x = 0^inf. write the equation first. y=lim (3x/1+2x2)^x then take the ln of both sides

OpenStudy (anonymous):

or is it because 0+ or 0- 0+ is inf 0- is -inf

OpenStudy (sirm3d):

1/0 is undefined. in limit problems, 1/0 = inf

OpenStudy (anonymous):

why do u still need ln it? dont we know that if we have 0^inf form, it must be 0 for sure.

OpenStudy (anonymous):

ooo

OpenStudy (sirm3d):

better safe than sorry.

OpenStudy (anonymous):

haha true, especially it is me who often make stupid mistakes

OpenStudy (anonymous):

last midterm, the ques asks for vertical asym, and i circled the one talking abt hori asym..>.>

OpenStudy (anonymous):

it's like i can never change this "careless" personality :(

OpenStudy (anonymous):

\[\frac{ +\infty }{ 0 }\] <what will this form look like?

OpenStudy (anonymous):

My note says "Very small # approaches -inf" , see if it approach from left or right"

OpenStudy (sirm3d):

inf (1/0) = inf (inf) = inf

OpenStudy (anonymous):

oh ok ,thanks. u are just too awesome :P

OpenStudy (anonymous):

I'm done my ques! will u be online anytime in the uncoming 26hrs to take last min ques @_@

OpenStudy (sirm3d):

i'll try.

OpenStudy (anonymous):

ok thanks!! but I think I just need to practice more and should be good thanks a lot alot :DDDDD

OpenStudy (anonymous):

Give all my hard ques to ur students and fail them ^^

OpenStudy (sirm3d):

i will. hahaha.

OpenStudy (sirm3d):

rest time...

OpenStudy (anonymous):

im gonna go coffee time :)

OpenStudy (anonymous):

see u later then :D

OpenStudy (anonymous):

@sirm3d you said inf / 0 = inf But in \[\frac{ e^x }{ \ln(1-\frac{ 2 }{ x } )} . \] The limit is - inf not +ve inf, why?

OpenStudy (anonymous):

How do I know which case in inf / 0 is +ve or -ve?

OpenStudy (sirm3d):

you only need to determine if it's 0+ or 0-, then follow the rule on signs. (-inf)/(0+)=(-/+) inf/0=(-) inf

OpenStudy (anonymous):

how do i know if it is 0+ or 0- ?

OpenStudy (anonymous):

Here i think is 0+ ?

OpenStudy (anonymous):